OpenStudy (anonymous):

what are the steps to solve: -9(-10+x)=126

OpenStudy (anonymous):

Divide both sides by negative nine. Add ten to both sides. The equation should then be in terms of x.

OpenStudy (anonymous):

could you show me how to solve the equation in a numerical form?

OpenStudy (anonymous):

-9(-10+x) / -9 =126 / -9 -10 + x = -14 -10 + x + 10 = -14 + 10 x = -4

OpenStudy (anonymous):

Thanks